1. An aqueous solution contains 7.65×10-2 M KHCO3 and 0.245 M H2CO3.  

The pH of this solution is

2. An aqueous solution contains 0.207 M KHCO3 and 0.466 M H2CO3.  

The pH of this solution is

3. A student measures the SO32- concentration in a saturated aqueous solution of silver sulfite to be 1.52×10-5 M.  
Based on her data, the solubility product constant for silver sulfite is

4. A student measures the molar solubility of zinc phosphate in a water solution to be 1.51×10-7 M.  
Based on her data, the solubility product constant for this compound is
Answer #1

1) Solution of KHCO3 and H2CO3 is a Buffer solution and it's pH is calculated by using Henderson's equation

pH = pKa1 + log [ Salt ] / [ Acid]

= 6.37 + log [KHCO3 ] / [H2CO3 ]

= 6.37 + log ( 7.65 x 10 -2) / 0.245

= 6.37 + log 0.312

= 6.37 - 0.506

= 5.86

2)

pH = pKa1 + log [ Salt ] / [ Acid]

= 6.37 + log [KHCO3 ] / [H2CO3 ]

= 6.37 + log ( 0.207) / 0.466

= 6.37 + log 0.444

= 6.37 - 0.352

= 6.02

3) Consider dissociation of Silver sulphite.

Ag2(SO3)(s) 2 Ag +(aq) + SO3 2-(aq)

[ SO3 2- ] = 1.52 x 10 -05 M then [ Ag + ] = 2 x 1.52 x 10 -05 M = 3.04 x 10 -05 M

We have , K sp = [Ag + ] 2 [ SO3 2-]  

= ( 3.04 x 10 -05 ) 2 x 1.52 x 10 -05

= 1.40 x 10 -14  

4)

Consider dissociation of Zinc phosphate.

Zn3(PO4)2(s) 3 Zn 2+(aq) +2 PO4 3-(aq)

Ksp = [Zn 2+ ] 3 [ PO4 3- ] 2

If S = Solubility of Zn3(PO4)2 in mol / L , then [ Zn 2+ ]  = 3 S and [PO4 3-] = 2 S

Therefore,   [ Zn 2+ ] = 3 x 1.51 x 10 -07 = 4.53 x 10 -07 mol / L

[PO4 3-] = 2 x 1.51 x 10 -07 =  3.02 x 10 -07 mol / L

Hence, Ksp = ( 4.53 x 10 -07) 3 x (  3.02 x 10 -07) 2

= 8.48 x 10 - 33